OpenStudy (anonymous):

"Determine whether each pair of lines is parallel, perpendicular, or neither. Q: 6x=5y+1 -12x+10y=1 (I know the answer is parallel but I need help working it out because I got the answer neither)

OpenStudy (texaschic101):

first put your equations in y = mx + b form where m is the slope and b is the y intercept. Then we will compare the slopes and the y intercepts. 6x = 5y + 1 6x - 1 = 5y -- divide both sides by 5 6/5x - 1/5 = y...lets rearrange y = 6/5x - 1/5 here the slope(m) is 6/5 and the y intercept (b) is -1/5 -12x + 10y = 1 -- add 12x to both sides 10y = 12x + 1 -- divide both sides by 10 y = 12/10x + 1/10 -- reduce y = 6/5x + 1/10 here the slope(m) is 6/5 and the y intercept(b) is 1/10 Now remember this.. if the slopes are the same, and the y intercepts are different, then it is a parallel line but if the slopes are the same and the y intercepts are the same, then it would be coincident...same line with infinite solutions. if the slopes are negative reciprocals of each other....that is if one slope is 2 and the other is 1/2, then the lines will be perpendicular. if the slopes are totally different...like 2 and 1/8...then it is neither

OpenStudy (anonymous):

Thank you so much! I have a huge Algebra test tomorrow and I didn't understand those. I have another but will work out by myself to see if I can get the answer on my own. I appreciate your help you rock :)